### v1

@@ Abstract
Lévy processes are widely used in financial modeling due to their ability to capture discontinuities and heavy tails, which are common in high-frequency asset return data. However, parameter estimation remains a challenge when associated likelihoods are unavailable or costly to compute. We propose a fast and accurate method for Lévy parameter estimation using the neural Bayes estimation (NBE) framework – a simulation-based, likelihood-free approach that leverages permutation-invariant neural networks to approximate Bayes estimators. Through extensive simulations across several Lévy models, we show that NBE outperforms traditional methods in both accuracy and runtime, while also enabling rapid bootstrap-based uncertainty quantification. We illustrate our approach on a challenging high-frequency cryptocurrency return dataset, where the method captures evolving parameter dynamics and delivers reliable and interpretable inference at a fraction of the computational cost of traditional methods. NBE provides a scalable and practical solution for inference in complex financial models, enabling parameter estimation and uncertainty quantification over an entire year of data in just seconds. We additionally investigate nearly a decade of high-frequency Bitcoin returns, requiring less than one minute to estimate parameters under the proposed approach.
@@ 1 Introduction
To better accommodate complex behavior such as discontinuities and heavy tails, financial researchers have long turned to Lévy processes – a flexible class of stochastic models characterized by stationary and independent increments. These processes are well established in financial econometrics and have been foundational for modeling returns for decades [Sato, 1999, Cont and Tankov, 2004, Applebaum, 2009, Bertoin, 1996, Madan et al., 1998].
Despite their flexibility, parameter estimation for Lévy processes remains a persistent challenge [Belomestny et al., 2015, Sueishi and Nishiyama, 2005, Kunitomo and Owada, 2006], primarily because many Lévy models lack tractable likelihoods. This is especially true for more expressive models, such as subordinated processes [Berry and Kleiber, 2023, Shirvani et al., 2024], which provide excellent fit in both central and tail regions but do not admit closed-form densities. As a result, classical likelihood-based methods are either computationally expensive or simply not applicable.
@@
Approximate Bayesian computation (ABC) has emerged as a promising alternative, particularly in combination with simulation-based strategies [Beaumont et al., 2002, Sisson et al., 2018, Chon and Cohen, 1997]. Recent developments in ABC have integrated neural networks to improve flexibility and scalability. Neural methods have been successfully applied across various domains, including genetics, ecology, econometrics, and spatio-temporal modeling [Creel, 2017, Flagel et al., 2018, Rudi et al., 2022, Gerber and Nychka, 2021, Banesh et al., 2021, Lenzi et al., 2023, Zammit-Mangion and Wikle, 2020]. These developments suggest that neural-based methods could offer a scalable alternative for Lévy parameter estimation.
In this work, we propose to use the neural Bayes estimation (NBE) framework [Sainsbury-Dale et al., 2023], a simulation-based approach that leverages neural networks to approximate Bayes estimators without requiring access to a likelihood. To implement this, we propose a neural architecture tailored for inference from independent replicates by leveraging permutation-invariant neural networks. Specifically, we build on the DeepSets framework [Zaheer et al., 2017], which provides an effective way to model functions of unordered independent and identically-distributed data.
@@
We evaluate this approach through several simulation studies involving common Lévy models and show that it consistently outperforms benchmark methods in both accuracy and computational efficiency. We then apply the method to high-frequency cryptocurrency return data.
Despite their clear utility for cryptocurrency modeling, Lévy processes have only recently been explored in this context. For instance, Shirvani et al. [2024] proposed a doubly subordinated normal inverse Gaussian Lévy process for daily Bitcoin return data, while Berry and Kleiber [2023] demonstrated that deeper subordination improves both tail modeling and small log-return behavior on high-frequency data. Although such models offer compelling fits, fast and scalable parameter estimation remains an open problem.
Our contributions in this work are threefold. First, we present a flexible, neural estimator for Lévy process parameters that bypasses the need for likelihood evaluation. Second, we implement the estimator using a DeepSets-based architecture, and investigate its performance in simple Lévy processes. Third, we demonstrate through a real-world cryptocurrency application that NBE delivers fast and accurate parameter estimation, outperforming traditional methods in both runtime and statistical precision. NBE enables us to estimate parameters and construct bootstrap-based confidence intervals for an entire year of high-frequency data in just seconds, marking a significant computational improvement over existing techniques.
@@ 2.1 Background on Lévy Processes
  • The distribution of $Z(t+h)-Z(t)$ does not depend on $t$.
@@
$\mathbb{E}[\text{exp}(\mathrm{i}\omega Z(t))]$ $=\text{exp}\Bigg{(}-\frac{t}{2}\sigma{\sigma}^{2}\omega^{2}+\mathrm{i}t{\gamma}\omega$
$\qquad+t\int\left[e^{\mathrm{i}\omega x}-1-\mathrm{i}\omega x\mathbbm{1}_{[|x|\leq 1]}(x)\right]{\nu}(dx)\Bigg{)}$
$=\text{exp}(t\Psi(w)),$ (1)
for $\omega\in\mathbb{R}$, where $\gamma\in\mathbb{R}$, $\sigma\geq 0$, and $\nu$ is a measure on $\mathbb{R}$ such that $\nu(\{0\})=0$ and $\int(|x|^{2}\wedge 1)\nu(dx)<\infty$
In this theorem, $\nu$ is the Lévy measure that governs the jump behavior of the process, $\gamma$ and $\sigma^{2}$ represent the drift and variance of the process, respectively, and $\Psi$ is the characteristic exponent. The triplet $\left(\gamma,\sigma,\nu\right)$ is referred to as the characteristic triplet of the Lévy process $Z(t)$ [Cont and Tankov, 2004], and fully defines the process structure.
@@ 2.1.2 Deep Variance Gamma Processes
A notable subclass of Lévy processes is that of subordinators: non-negative, almost surely increasing processes in time. The key property of subordinators is that if $Z(t)$ is a Lévy process and $S(t)$ is a subordinator, then the time-changed process $Z(S(t))$ is also a Lévy process. This property allows subordinators to serve as building blocks for constructing more flexible Lévy processes via random time changes (see, for example, the VG process defined in Section 2.1.1).
@@
As shown in Berry and Kleiber [2023], the DVG model provides a strong framework for capturing the empirical behavior of higih-frequency cryptocurrency datasets. However, for orders $L\geq 2$, the model lacks a closed-form likelihood expression, which presents challenges for standard likelihood-based inference methods.
@@ 2.2 Estimation Methods
We now turn to the problem of estimating parameters from discretely-observed Lévy processes. This topic has been widely studied, with most traditional methods relying on the likelihood function [Cont and Tankov, 2004, Sueishi and Nishiyama, 2005, Barndorff-Nielsen and Blaesild, 1981]. However, for many Lévy models, including the DVG process with depth $L\geq 2$, the likelihood function cannot be expressed in closed form [Diggle and Gratton, 1984, Berry and Kleiber, 2023]. In these situations, alternative approaches such as maximum empirical likelihood [Elgin, 2011] and nonparametric estimation techniques [Figueroa-López, 2009] have been proposed. We focus on two such approaches, assuming we observe data $\mathcal{X}=\{X_{j}\}_{j=1}^{n}$ generated from a Lévy process.
We begin with the least squares characteristic function (LSQ) approach. This method estimates parameters by minimizing the squared average difference between the empirical and theoretical characteristic functions across a selected set of frequencies [Yu, 2004, Xu and Darve, 2020, Kappus and Reiß, 2010]. This method was applied successfully to the DVG process in [Berry and Kleiber, 2023], where it showed promising results in terms of accuracy.
Let $\phi(\omega,\theta)$ denote the theoretical characteristic function of a Lévy process with parameter vector $\theta$, evaluated at frequency $\omega\in\mathbb{R}$. The empirical characteristic function, $\phi_{n}(\omega)$, is computed from the dataset as $\phi_{n}(\omega)=\frac{1}{n-1}\sum_{j=1}^{n-1}\text{exp}(i\omega\widetilde{X_{j}})$. Where $\widetilde{X_{j}}=X_{j+1}-X_{j}$ for $j=1,\ldots,n-1$. The LSQ method then estimates $\hat{\theta}$ by minimizing the discrepancy between the empirical and theoretical characteristic functions,
@@
A second widely-used method is maximum empirical likelihood estimation (MELE) [Qin and Lawless, 1994], which also makes use of the empirical characteristic function. The core idea behind MELE is to assign a probability $p_{j}$ to each observed data point $\widetilde{X_{j}}$, and then find the set of probabilities that best matches a set of moment conditions derived from the characteristic function.
Formally, we define the empirical likelihood function as $L(F_{\theta})=\prod p_{j}$, where $F_{\theta}$ represents the distribution function associated with parameter $\theta$, and the $p_{j}$ are subject to constraints that ensure consistency with the empirical data. Consider the empirical distribution function $F_{n}(x)=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}_{[X_{i}<x]}$ where $\mathbbm{1}$ is the indicator function that is unity when the set condition holds, zero otherwise. Using this, the empirical likelihood ratio is defined as $R(F)=L(F)/L(F_{n})$, where the denominator corresponds to the empirical distribution with equal weights. MELE seeks to maximize this ratio subject to the following constraints:
$p_{i}\geq 0,\quad\sum_{i=1}^{n}p_{i}=1,\quad\sum_{i=1}^{n}p_{i}[e^{i\omega X_{i}}-\phi(\omega,\theta_{L})]=0.$
In essence, MELE chooses the probabilities $p_{i}$ such that the weighted average of the empirical characteristic function matches the theoretical characteristic function implied by $\theta$.
@@ 2.2.1 Neural Bayes Estimation
One particularly attractive property of Lévy processes is that they are often straightforward to simulate, lending themselves well to simulation-based estimation techniques. Neural networks have recently emerged as a powerful tool for likelihood-free inference. In particular, Sainsbury-Dale et al. [2023] demonstrated the effectiveness of neural point estimators for parameter estimation from replicated samples, using different spatial models such as Gaussian processes, max-stable processes, and conditional extremes models.
@@
[…]
@@
Note that this approach does not require evaluation or explicit knowledge of the likelihood function. After training, we refer to the fitted neural point estimator as a neural Bayes estimator. Although training the neural network can be computationally expensive, the cost is amortized over time. Once trained, the model can generate parameter estimates for new datasets with minimal computational overhead while also providing fast, bootstrap-based uncertainty quantification.
2.2.2 DeepSets
A unique aspect of Lévy processes that we wish to exploit within this architecture is the stationary and independent increments. That is, the distribution of the vector $\widetilde{\mathbf{X}}=(\widetilde{X}_{1},\widetilde{X}_{2},\ldots,\widetilde{X}_{n-1})^{\rm T}=(X_{2}-X_{1},X_{3}-X_{2},\ldots,X_{n}-X_{n-1})^{\rm T}$ is invariant to permutations. Thus it is desirable to use a neural network architecture that exploits this invariance. We follow Sainsbury-Dale et al. [2023], which is based on the DeepSets framework [Zaheer et al., 2017]. DeepSets enforces the permutation invariance property by structuring the neural estimator as follows:

### v2

@@ Abstract
Lévy processes are widely used in financial modeling due to their ability to capture discontinuities and heavy tails, which are common in high-frequency asset return data. However, parameter estimation remains a challenge when associated likelihoods are unavailable or costly to compute. We propose a fast and accurate method for Lévy parameter estimation using the neural Bayes estimation (NBE) framework – a simulation-based, likelihood-free approach that leverages permutation-invariant neural networks to approximate Bayes estimators. We contribute new theoretical results, showing that NBE results in consistent estimators whose risk converges to the Bayes estimator under mild conditions. Moreover, through extensive simulations across several Lévy models, we show that NBE outperforms traditional methods in both accuracy and runtime, while also enabling two complementary approaches to uncertainty quantification. We illustrate our approach on a challenging high-frequency cryptocurrency return dataset, where the method captures evolving parameter dynamics and delivers reliable and interpretable inference at a fraction of the computational cost of traditional methods. NBE provides a scalable and practical solution for inference in complex financial models, enabling parameter estimation and uncertainty quantification over an entire year of data in just seconds. We additionally investigate nearly a decade of high-frequency Bitcoin returns, requiring less than one minute to estimate parameters under the proposed approach.
@@ 1 Introduction
To better accommodate complex behavior such as discontinuities and heavy tails, financial researchers have long turned to Lévy processes – a flexible class of stochastic models characterized by stationary and independent increments. These processes are well established in financial econometrics and have been foundational for modeling returns for decades [Bertoin, 1996, Madan et al., 1998, Sato, 1999, Cont and Tankov, 2004, Applebaum, 2009].
Despite their clear utility for cryptocurrency modeling, Lévy processes have only recently been explored in this context. For instance, Shirvani et al. [2024] proposed a doubly subordinated normal inverse Gaussian Lévy process for daily Bitcoin return data, while Berry and Kleiber [2023] demonstrated that deeper subordination improves both tail modeling and small log-return behavior on high-frequency data.
Although such models offer compelling fits, fast and scalable parameter estimation remains a persistent challenge [Sueishi and Nishiyama, 2005, Kunitomo and Owada, 2006, Belomestny et al., 2015], primarily because many Lévy models lack tractable likelihoods. This is especially true for more expressive models, such as subordinated processes [Berry and Kleiber, 2023, Shirvani et al., 2024], which have been shown to capture both central and tail behavior effectively in many financial datasets, but do not admit closed-form densities. As a result, classical likelihood-based methods are either computationally expensive or simply not applicable.
@@
Given these limitations, simulation-based Bayesian approaches such as approximate Bayesian computation (ABC) have become an attractive alternative [Chon and Cohen, 1997, Beaumont et al., 2002, Sisson et al., 2018]. Recent developments in ABC have integrated neural networks to improve flexibility and scalability. In particular, there has been substantial work on neural point estimators: neural networks designed to learn a direct mapping from the observed data to the parameter space, producing point estimates rather than full posterior distributions. Such neural methods have been successfully applied across various domains, including genetics, ecology, econometrics, and spatio-temporal modeling [Creel, 2017, Flagel et al., 2018, Zammit-Mangion and Wikle, 2020, Gerber and Nychka, 2021, Banesh et al., 2021, Rudi et al., 2022, Lenzi et al., 2023]. These developments suggest that neural-based methods could offer a scalable alternative for Lévy parameter estimation.
Neural networks have also gained significant attention within the financial community, where they are valued for the improvements in speed and accuracy they can provide. Substantial work has focused on their application to model calibration and option pricing. For instance, Hernandez [2016] demonstrated that neural networks can be used effectively for model calibration in settings where the likelihood is intractable. Subsequent works, such as Liu et al. [2019], Gan et al. [2020], further developed neural frameworks for model calibration. Horvath et al. [2021] introduced a method in which neural networks approximate complex pricing functions that are otherwise difficult to evaluate. See [Ruf and Wang, 2020] for a review of this growing literature on neural networks for option pricing and hedging.
More recently, neural networks have also been applied directly in the context of Lévy processes within the financial literature. For example, Huh [2019] proposed a new non-parametric model, the exponential Lévy neural network (ELNN), for option pricing. Kudryavtsev and Rodochenko [2020], Kudryavtsev and Danilova [2023] employed neural networks to simulate and calibrate different classes of Lévy processes. These works underscore a growing trend at the intersection of machine learning and Lévy process modeling.
In this work, we propose to use the neural Bayes estimation (NBE) framework [Sainsbury-Dale et al., 2023], a simulation-based approach that leverages neural networks to approximate Bayes estimators without requiring access to a likelihood. Our contributions are fourfold. First, we introduce a flexible neural Bayes estimator for Lévy process parameters that bypasses the need for likelihood evaluation. Second, we implement the estimator using a DeepSets-based architecture [Zaheer et al., 2017], which is tailored for inference from replicates of unordered i.i.d. data. Third, we provide new theoretical results showing that the estimator is consistent and converges in risk to the Bayes estimator, offering formal guarantees on its large-sample behavior. Fourth, we extend uncertainty quantification by comparing two complementary approaches — bootstrap-based confidence intervals and posterior-quantile estimation — thereby providing a principled and interpretable framework for likelihood-free inference.
@@
We evaluate this approach through several simulation studies involving common Lévy models and demonstrate that it consistently outperforms benchmark methods in both accuracy and computational efficiency. Finally, we apply the method to high-frequency cryptocurrency return data, where NBE delivers fast and accurate parameter estimation, outperforming traditional methods in both runtime and statistical precision. Remarkably, our approach allows us to estimate parameters and construct confidence intervals for an entire year of high-frequency data in just seconds, marking a significant computational improvement over existing techniques.
@@ 2.1 Background on Lévy Processes
  • The distribution of Z(t+h)Z(t)Z(t+h)-Z(t) does not depend on $t$.
@@
$\mathbb{E}[\exp(\mathrm{i}\omega Z(t))]$ $=\exp\Bigg(-\frac{t}{2}{\sigma}^{2}\omega^{2}+\mathrm{i}t{\gamma}\omega$
$\hskip 18.49988pt+t\int\left[e^{\mathrm{i}\omega x}-1-\mathrm{i}\omega x\mathbbm{1}_{[|x|\leq 1]}(x)\right]{\nu}(dx)\Bigg)$
$=\exp(t\Psi(w)),$ (1)
for $\omega\in\mathbb{R}$, where $\gamma\in\mathbb{R}$, $\sigma\geq 0$, and $\nu$ is a measure on $\mathbb{R}$ such that $\nu(\{0\})=0$ and $\int(|x|^{2}\wedge 1)\nu(dx)<\infty$ .
In this theorem, $\mathbbm{1}$ denotes the indicator function, $\nu$ is the Lévy measure that governs the jump behavior of the process, $\gamma$ and $\sigma^{2}$ represent the drift and variance of the process, respectively, and $\Psi$ is the characteristic exponent. The triplet $\left(\gamma,\sigma,\nu\right)$ is referred to as the characteristic triplet of the Lévy process $Z(t)$ [Cont and Tankov, 2004], and fully defines the process structure.
@@ 2.1.2 Deep Variance Gamma Processes
A notable subclass of Lévy processes is that of subordinators: non-negative, almost surely increasing processes in time [Cont and Tankov, 2004]. The key property of subordinators is that if $Z(t)$ is a Lévy process and $S(t)$ is a subordinator, then the time-changed process $Z(S(t))$ is also a Lévy process. This property allows subordinators to serve as building blocks for constructing more flexible Lévy processes via random time changes (see, for example, the VG process defined in Section 2.1.1).
@@
As shown in Berry and Kleiber [2023], the DVG model provides a strong framework for capturing the empirical behavior of high-frequency cryptocurrency datasets. However, for orders $L\geq 2$, the model lacks a closed-form likelihood expression, which presents challenges for standard likelihood-based inference methods.
@@ 2.2 Estimation Methods
We now turn to the problem of estimating parameters from discretely-observed Lévy processes. This topic has been widely studied, with most traditional methods relying on the likelihood function [Barndorff-Nielsen and Blaesild, 1981, Cont and Tankov, 2004, Sueishi and Nishiyama, 2005]. However, for many Lévy models, including the DVG process with depth $L\geq 2$, the likelihood function cannot be expressed in closed form [Diggle and Gratton, 1984, Berry and Kleiber, 2023]. In these situations, alternative approaches such as maximum empirical likelihood [Elgin, 2011] and nonparametric estimation techniques [Figueroa-López, 2009] have been proposed. We focus on two such classical approaches, namely the least squares characteristic function (LSQ) and maximum empirical likelihood (MELE) methods. In addition, we introduce Neural Bayes Estimators (NBEs) and describe how they can be applied to parameter estimation for Lévy processes.
2.2.1 Classical Methods
Assume we observe data $\mathcal{X}=\{X_{j}\}_{j=1}^{n}$ generated from a Lévy process, we begin with the LSQ characteristic function approach. This method estimates parameters by minimizing the squared average difference between the empirical and theoretical characteristic functions across a selected set of frequencies [Yu, 2004, Kappus and Reiß, 2010, Xu and Darve, 2020]. This method was applied successfully to the DVG process in [Berry and Kleiber, 2023], where it showed promising results in terms of accuracy.
[…]
@@ 2.2.1 Classical Methods
A second widely used method is MELE [Qin and Lawless, 1994], which also makes use of the empirical characteristic function. The core idea behind MELE is to assign a probability $p_{j}$ to each observed data point $\widetilde{X_{j}}$, and then find the set of probabilities that best matches a set of moment conditions derived from the characteristic function.
Formally, MELE solves the following constrained optimization problem:
$\max_{p_{1},\dots,p_{n}}\;\sum_{i=1}^{n}\log p_{i}$
subject to
$p_{i}\geq 0,\qquad\sum_{i=1}^{n}p_{i}=1,\qquad\sum_{i=1}^{n}p_{i}\,g(X_{i},\theta)=0,$
where $g(X_{i},\theta)$ represent the moment conditions derived from the model. In the Lévy setting, these conditions are given by
$g(X_{i},\theta)=e^{i\omega X_{i}}-\phi(\omega,\theta).$
In other words, MELE chooses the probabilities $p_{i}$ so that the weighted empirical characteristic function matches the theoretical characteristic function implied by $\theta$.
@@ 2.2.2 Neural Bayes Estimation
One particularly attractive property of Lévy processes is that they are often straightforward to simulate, lending themselves well to simulation-based estimation techniques. Neural networks have recently emerged as a powerful tool for likelihood-free inference. Sainsbury-Dale et al. [2023] demonstrated the effectiveness of neural point estimators for parameter estimation from replicated samples, using different spatial models such as Gaussian processes, max-stable processes, and conditional extremes models.
@@
$R_{\Omega}\left(\hat{\theta}(\cdot)\right)=\int_{\Theta}\left(\int_{\mathcal{S}}L\left(\theta,\hat{\theta}(\mathcal{X})\right)f\left(x|\theta\right)dx\right)d\Omega(\theta)$ (2)
@@
Note that this approach does not require evaluation or explicit knowledge of the likelihood function. After training, we refer to the fitted neural point estimator as a neural Bayes estimator. Although training the neural network can be computationally expensive, the cost is amortized over time – that is, there is an upfront cost in training the network, but executing the fitted network on new datasets is very fast. Once trained, the model can generate parameter estimates for new datasets with minimal computational overhead while also providing fast uncertainty quantification.
2.2.3 Loss Functions
It is well known that in Bayesian decision theory, the choice of a loss function determines which functional of the posterior distribution is approximated by the Bayes estimator Lehmann and Casella [1998]. For example, under squared error loss, $L(\theta,\hat{\theta})=\|\theta-\hat{\theta}\|^{2}$, a simple derivation shows that the risk is minimized when the estimator is the posterior mean. Thus, the squared error loss is said to be paired with the posterior mean.
This observation is particularly relevant in the context of neural Bayes estimators, since the loss function chosen during training directly determines which posterior functional the network learns to approximate. For example, using the asymmetric lin-lin loss with different $\alpha$ values allows NBE to approximate posterior $\alpha$-quantiles, and thereby construct likelihood-free credible intervals.
2.2.4 Uncertainty Quantification
Once the neural Bayes estimator is trained, uncertainty quantification (UQ) can be performed in multiple ways. We consider two complementary approaches below.
Following Sainsbury-Dale et al. [2023], uncertainty can be quantified within the NBE framework using a nonparametric bootstrap procedure. In this approach, multiple bootstrap datasets are generated by resampling (with replacement) from the observed data, and the neural Bayes estimator is reevaluated for each dataset. Confidence intervals can then be obtained from the empirical distribution of these estimates. Such intervals represent uncertainty in the NBE procedure itself: if $\hat{\theta}(\mathcal{X})$ denotes the estimator trained on data $\mathcal{X}$, bootstrap-based variance (for example) approximates $\mathrm{Var}\big(\hat{\theta}(\mathcal{X})\big)$ capturing the sampling variability of the estimator, as opposed to other aspects of the posterior distribution.
Alternatively, uncertainty can be quantified by directly interrogating aspects of the posterior distribution, leveraging the connection between loss functions and Bayes estimators. For example, training the network with the lin-lin loss
$L(\theta,a)=\big(I\{\theta\leq a\}-\alpha\big)(a-\theta),\qquad\alpha\in(0,1),$
yields the $\alpha$-quantile of the posterior distribution. By training separate networks for different values of $\alpha$, one can recover posterior credible intervals that reflect the Bayesian parameter uncertainty captured by the posterior itself.
A trade-off between these approaches lies in computational cost. Estimating multiple posterior quantiles can be computationally intensive, whereas bootstrap-based UQ is generally faster, especially when simulations are readily available. However, bootstrap only captures the variability of the estimator itself, rather than the full posterior uncertainty. For a more detailed discussion on estimating the posterior distribution, or guidance on when each method may be preferred, we refer the reader to Zammit-Mangion et al. [2025], and Sainsbury-Dale et al. [2025].
2.2.5 DeepSets Architecture for Lévy Processes
Having defined the NBE framework and the role of the loss function, we now describe the specific neural network architecture used in this work. Here, architecture refers to the actual structure of the neural network that determines how inputs are mapped to outputs.
A unique aspect of Lévy processes that we wish to exploit is their stationary and independent increments. That is, the distribution of the vector $\widetilde{\mathbf{X}}=(\widetilde{X}_{1},\widetilde{X}_{2},\ldots,\widetilde{X}_{n-1})^{\rm T}=(X_{2}-X_{1},X_{3}-X_{2},\ldots,X_{n}-X_{n-1})^{\rm T}$ is invariant to permutations. Thus it is desirable to use a neural network architecture that exploits this invariance. We follow Sainsbury-Dale et al. [2023], which is based on the DeepSets framework [Zaheer et al., 2017]. DeepSets enforces the permutation invariance property by structuring the neural estimator as follows: